a ball whose mass is .35 kg falls from rest through a height of 5.3 m and then bounces back. it reaches a maximum height of 3.8 m after the bounce before falling to the ground again. from the beginning at 5.3 m to the height of 3.8 m after the bounce. the magnitude of its mechanical energy has changed by

a) zero

b) 13.0 J

c) 0.52 J

d) 5.15 J
Answer #1

We know that Mechanical energy is given by sum of kinetic energy and Potential energy.

At height 5.3 meter Kinetic energy is zero at that time and potential energy is M*g*h1 =.35kg×9.8(m/sec^2)×5.3m=18.179 joule = Intial Mechanical energy

After bouncing back to height 3.8m Kinetic energy is zero at that time and potential energy is given by M*g*h2=0.35kg×9.8(m/sec^2)×3.8=13.034 joule = final mechanical energy

Therefore change in mechanical energy is given by

( Intial Mechanical energy-Final mechanical energy)

=18.149-13.034

= 5.115 joule

Hence option D is correct

Note - here I use the Standard value of g =9.8m/sec^2